Dynamic Launch Decisions for Satellite Constellation

Operators

William King

Washington State University

November 8, 2020

Abstract

Over the last 10 years new technology has make low earth orbits (LEOs) more accessible, and the resulting increase in LEO satellites has increased the risk of collision. Because debris in orbit generates more debris through collisions with objects in orbit and the debris created during launch and operation imposes a negative externality on other operators, optimal use of orbits is believed to not occur under free entry. This paper develops a dynamic model of satellite operation incorporating two effects not considered in previous models. The first effect is complementarity between satellites within the same operator’s fleet (called a constellation). The second effect is collision avoidance efficiencies that exist within constellations. The primary result is a theoretical model and the resulting analysis of the difference in survival ratios between constellation operators and society.

Keywords: Orbits, Pollution, Economies of Scale, Externality

JEL Codes: Q29, Q58, L25 Acknowledgments: I am the sole author and have recieved no contributions from others as of yet. This paper has been approved for dual submission in Econs 529 and Econs 594 by the instructors.